OpenStudy (anonymous):

A newly discovered planet has a mass 1.5 times compared to that of the earth, and its diameter is 3 times compared to earth. What is the gravitational acceleration on this planet?

OpenStudy (irishboy123):

we relate the idea of g at the earth's surface to Newton's law of universal gravitation, such that at the surface of our planet of radius \(R_e\), we can say: \(\large F = \frac{GM_em}{R_e^2} = mg_e\) so \(\large g_e = \frac{GM_e}{R_e^2}\) Then, noting that G is a universal constant, we can say that: \(\large G = \frac{g_e \ R_e^2}{M_e}\) such that, as between 2 different planets, earth and planet X: \( \huge \frac{g_e \ R_e^2}{M_e} = \frac{g_x \ R_x^2}{M_x} \) and so, to find \(g_x\) using the data provided in the question

OpenStudy (anonymous):

what would be my Re and Me?

OpenStudy (anonymous):

is it 8.42 x10 ^-23?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

is it 3.021 x 10-34?

OpenStudy (irishboy123):

i was teasing. my apologies. we know from above equations that: \(\large g_x = g_e \times \frac{R_e^2}{Me} \times \frac{M_x}{R_x^2} = g_e \times (\frac{R_e}{R_x})^2 \times \frac{M_x}{M_e}\) you have everything [ie the ratios!!] you need in the original question to do this.

OpenStudy (irishboy123):

from the question: .....**has a mass 1.5 times compared to that of the earth***..... ***diameter is 3 times compared to earth**** so you know \(\frac{M_x}{M_e}\) and \(\frac{R_x}{R_e}\)